OpenStudy (anonymous):

Multiplication trick of 11 by Muhammad: Hi friends this is trick with you can multiply by 11 in seconds Example: 24*11= Add 2+4=6 then insert 6 between in 2 and 4=264 Another example: 55*1=605 1st step: Add both numbers 5+5=10 then insert 10 between 55 in this we have add 5+105 605 answer. You can check by using calculator.. Comments please.

OpenStudy (anonymous):

anyway nice :)

OpenStudy (lgbasallote):

the 2nd example is vague

OpenStudy (lgbasallote):

also how would you use it in something like 100 * 11

OpenStudy (anonymous):

nys..

OpenStudy (anonymous):

thanks for sharing, there are many methods out there. here is one i like to share 12 x 34 ------ 1. multiply starting from right vertical 2*4 = 8 12 x 34 ------ 8 2. cross multiply top digit and bottom digit and add together " carry a digit if need" (1* 4 ) + (2*3) = 10 1 <-------- the carry digit 12 x 34 ------ 0 8 3. multiply the the vertical left. " add the carry digit if there is one" (3*1 )+1 = 4 1 <-------- the carry digit 12 x 34 ------ 4 0 8 this work with all 2 digit multiplication